hey,
i have a 2 word problems and could you please solve it for me, and thanks.
1. find two consecutive integers with a product 56.
2. The sum of the squares of two consecutive integers is 145. find the integers.
i will be waiting for the solutions and thank you again.

Answer by checkley77(12844)

X(X+1)=56
X^2+X-56=0
(X+8)(X-7)=0
X=7 FOR THE FIRST INTEGER
8+1=9 FOR THE SECOND.
PROOF:
7*8=56
56=56
OR:
x=8=0
x=-8 answer.
-8+1=-7 answer.
--------------------------
X^2+(X+1)^2=145
X^2+X^2+2X+1-145=0
2X^2+2X-144=0
2(X^2+X-72)=0
2(X+9)(X-8)=0
X-8=0
X=8 FOR THE SMALLER INTEGER.
8+1=9 FOR THE LARGER INTEGER.
PROOF:
8^2+9^2=145
64+81=145
145=145
OR:
x+9=0
x=-9 answer.
-9+1=-8 answer.